Listed below are speeds​ (mi/h) measured from traffic on a busy highway. This simple random sample was obtained at​ 3:30 P.M. on a weekday. Use the sample data to construct a 90​% confidence interval estimate of the population standard deviation. 60,65,65,55,65,53,59,58,59,70,61,67
Answer #1

(n-1)s*

s^2 = 25.53788

df = n-1 =11

sqrt(25.5378 * 11 / 19.67514 ) < sd < sqrt( 25.5378 * 11 / 4.574813)

= (3.778585, 7.836126)